Part A econd and third dark lines of the force there when their there w eg
Answer #1

expression for fringe width

x = (n3 - n2) L D / d

x = ( 3 - 2) * 540* 10^-9* 0.725 / 0.45

x = 8.7* 10^-7 m

======

Comment before rate in case any doubt, will reply for sure.. goodluck